If you enclose a rectangular garden using a side of a building as one side of the rectangle, what are the dimensions of the garden if it is to be the maximum area that you can enclose with 42 feet of fence?
Answer by stanbon(75887)

If you enclose a rectangular garden using a side of a building as one side of the rectangle, what are the dimensions of the garden if it is to be the maximum area that you can enclose with 42 feet of fence?
----------
Draw the picture.
The two equal sides (perpendicular to the bldg.) are "x" long.
The side parallel to te bldg is 42-2x long
-------------
Equation:
A(x) = x(42-2x) = 42x - 2x^2
------
You have a quadratic with a = -2 ; b = 42
----
Max area occurs where x = -b/(2a) = -42/(2*-2) = 42/4 = 10.5 feet
=========================
42-2x = 42-21 = 21 feet
==========================
Cheers,
Stan H.